OpenStudy (anonymous):

Can someone please help? :) List three different ways to write 5^11 as the product of two powers. Explain why all three of your expressions are equal to 5^11

OpenStudy (anonymous):

5^11 or 2^5+2^5+2^1, cause it all adds up or 4^10+1^1

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (jdoe0001):

2^5+2^5 = 64 2^5*2^5 = 1024 $$ 2^5+2^5 \ne 2^5\times 2^5 $$